# Understanding Impact of Human Feedback via Influence Functions

### v1

@@ B.2 Labeling strategy oversight
Dataset Source Label Accuracy Size (Train) Size (Validation)
Helpsteer2 (Train) 72.96$\pm$1.04 8218 432
Table 4: Label Accuracy denotes the proportion of cases where Bob’s preference labels match those of Alice in $\mathcal{D}_{\tt B}$.
@@
Appendix C Baselines
C.1 LLM baselines
In the length bias experiment, the 3-shot examples consisted of conversations where the chosen response was deliberately more concise, depending on the prompt structure. The model was tasked with evaluating both the brevity and the quality of the content. Additionally, in the sycophancy bias experiment, the 3-shot examples included scenarios where the chosen response was factually accurate but less aligned with the user’s opinion, while the rejected response exhibited sycophantic tendencies. The prompts used for 3-shot learning in these LLM experiments can be found in  subsection C.1, and the data used can be found in  Figure 9 and  Figure 10
@@ C.1 LLM baselines
Table 5: Prompt format used for few-shot prompting.
Figure 9: Prompt and response pair used for few-shot prompting in length bias experiments, including chosen and rejected responses. Samples without length bias were intentionally selected for the experiment.
Figure 10: Prompt and response pair used for few-shot prompting in sycophancy bias experiments, including chosen and rejected responses. Samples without sycophancy bias were intentionally selected for the experiment.
@@ C.2.3 Self-Confidence and Entropy
We also adopt two additional baselines for bias detection experiments that evaluate label quality based on training data: self-confidence and entropy. Both are derived from the model’s predicted probabilities for the winning response $y^{(z)}$ and the losing response $y^{(1-z)}$. To maintain consistency with the influence and Mahalanobis distance metrics—where higher values indicate more biased behavior—we reversed the signs of both the self-confidence and entropy metrics, ensuring that higher values for these metrics also reflect greater bias.
@@
where $p_{y^{(z)}}$ is the predicted probability of the winning response, derived from the softmax transformation of the logits. Originally, a lower self-confidence score indicated more biased behavior, but with the sign reversal, a higher self-confidence score now reflects greater bias in the model’s predictions.
@@
$\text{Entropy}=\sum_{z\in\{0,1\}}p_{y^{(z)}}\log(p_{y^{(z)}})$
where $p_{y^{z}}$ represents the probability for response ($y^{(0)}$ or $y^{(1)}$). Initially, lower entropy indicates greater confidence in one response and thus less bias. With the sign reversed, higher entropy values now indicate greater uncertainty and, consequently, greater bias.
@@ 3. The output should be a JSON object of the following format:
## Warning: please only focus on the ”Response to evaluate” part when you measure the sycophancy based on the score rubrics.
@@ Appendix E Additional metrics for bias detection
Bias Type Length Sycophancy
AUC AP TNR80 AUC AP TNR80
Influence 0.800 0.237 0.664 0.711 0.121 0.489
Confidence 0.616 0.094 0.361 0.585 0.064 0.297
Entropy 0.589 0.079 0.333 0.533 0.045 0.278
Mahalanobis 0.576 0.082 0.277 0.560 0.052 0.237
KNN 0.582 0.083 0.303 0.533 0.047 0.230
@@ Appendix F Alice and bob experiment weight update method
Influence-based partitioning
Training the SVM classifier
Cosine similarity and accuracy evaluation
@@
Appendix G Qualitative analysis
@@
To investigate length bias, we used the Concise dataset, focusing on cases where response length may affect outcomes. As shown in Figure 12, the most positively-contributing samples, which negatively impact bias, show no flipped instances, and the selected responses demonstrate no preference for length. This indicates that these responses did not favor verbose answers.
@@
For sycophancy bias, we examined the Less Sycophantic dataset. As illustrated in Figure 13, most positively-contributing samples do not contain any flipped instances, and none of the responses show sycophancy bias. This implies that these responses remained neutral and did not overly agree with the user’s views.
Figure 12: The most positively-contributing samples do not contain any flipped instances, and the selected responses show no evidence of length bias. In contrast, most negatively-contributing samples include three out of four instances that were flipped, with the remaining instance also displaying length bias.
Figure 13: The most positively-contributing samples do not include any flipped instances, and the selected responses show no signs of sycophancy bias. In contrast, most negatively-contributing samples include one flipped instance, with all exhibiting sycophancy bias.
@@
Appendix H Ablation experiments on bias detection
H.1 Validation set size ablation for influence functions
@@
The ablation results of the validation set size are given in Figure 14 and Figure 15. These results demonstrate that influence functions are capable of accurately detecting both biases with as few as 50 samples. Furthermore, the performance reaches saturation after 50 samples for length bias, and 100 samples for sycophancy bias, indicating that increasing the validation set size beyond this point yields diminishing returns. This efficiency suggests that influence functions can effectively capture critical patterns in the preference dataset, even when using a relatively small validation set of 50 samples.
H.2 Few-shot example ablation for llm baselines

### v2

@@ B.2 Labeling Strategy Oversight
Dataset Source: Helpsteer2 (Train)
Label Accuracy: $72.96\pm 1.04$
Size (Train): 8218
Size (Validation): 432
Table 5: Details of the dataset used in the labeling strategy oversight experiment. Label Accuracy denotes the proportion of cases where Bob’s preference labels match those of Alice in $\mathcal{D}_{\tt B}$.
@@
Table 6: Prompt format used for few-shot prompting LLM baselines in bias detection experiments.
Appendix C Baselines
C.1 LLM Baselines
In the length bias experiment, the 3-shot examples consisted of conversations where the chosen response was deliberately more concise, depending on the prompt structure. The model was tasked with evaluating both the brevity and the quality of the content. Additionally, in the sycophancy bias experiment, the 3-shot examples included scenarios where the chosen response was factually accurate but less aligned with the user’s opinion, while the rejected response exhibited sycophantic tendencies. The prompts used for 3-shot learning in these LLM experiments can be found in  subsection B.2, and the few shot examples used can be found in  Figure 6 and  Figure 7
Figure 6: Prompt and response pair used for few-shot prompting of LLM baselines in length bias experiments, including chosen and rejected responses. Samples without length bias were intentionally selected for the experiment.
Figure 7: Prompt and response pair used for few-shot prompting for LLM baselines in sycophancy bias experiments, including chosen and rejected responses. Samples without sycophancy bias were intentionally selected for the experiment.
@@ C.2.3 Self-Confidence and Entropy
We also adopt two additional baselines for bias detection experiments that evaluate label quality based on training data: self-confidence and entropy. Both are derived from the model’s predicted probabilities for the winning response $y^{(z)}$ and the losing response $y^{(1-z)}$. To maintain consistency with the influence and Mahalanobis distance metrics, where higher values indicate more biased behavior, we reversed the signs of both the self-confidence and entropy metrics, ensuring that higher values for these metrics also reflect greater bias.
@@
where $p_{y^{(z)}}$ is the predicted probability derived from the softmax transformation of the logits.
@@
$\text{Entropy}=-\sum_{z\in\{0,1\}}p_{y^{(z)}}\log(p_{y^{(z)}})$
where $p_{y^{(z)}}$ represents the probability for response ($y^{(0)}$ or $y^{(1)}$).
@@ 3. The output should be a JSON object of the following format:
## Warning: please only focus on the "Response to evaluate" part when you measure the sycophancy based on the score rubrics.
@@ Appendix E Additional Metrics for Bias Detection
Length Bias
Bias Type AUC AP TNR80
Influence 0.800 0.237 0.664
Confidence 0.616 0.094 0.361
Entropy 0.589 0.079 0.333
Mahalanobis 0.576 0.082 0.277
KNN 0.582 0.083 0.303
Sycophancy Bias
Bias Type AUC AP TNR80
Influence 0.711 0.121 0.489
Confidence 0.585 0.064 0.297
Entropy 0.533 0.045 0.278
Mahalanobis 0.560 0.052 0.237
KNN 0.533 0.047 0.230
@@ Appendix F Alice and Bob Experiment Weight Update Method
Influence-Based Partitioning
Training the SVM Classifier
Cosine Similarity and Accuracy Evaluation
Appendix G Reward Retraining on Curated Set Using Bias Detection
In this section, we evaluate whether flipping the preference labels of negatively-contributing samples detected by influence functions improves reward model performance, measured by validation set accuracy. We compare retrained models against reward models trained on the datasets from our main analysis (referred to as main datasets). Our main datasets contain manually flipped preference labels. We also compare them with a clean dataset without any label modifications. Specifically, we flip $\alpha$-($\alpha={50,100,150}$) of the actual number of manually flipped preference labels from the training dataset. For example, for $\alpha=50$% we flip the top 3.28% samples with the highest influence scores as the length training dataset contains 6.56% of manually flipped preference labels. We retrain the reward model using this curated dataset and then report its validation set accuracy. For length bias experiments, Concise and Verbose validation set accuracies are reported. For sycophancy bias experiments, Less Sycophantic and More Sycophantic validation set accuracies are reported. Please refer to Appendix B for details on these validation sets.
As shown in Table 10, flipping preference labels of negatively-contributing samples increase Concise validation accuracy up to 14% and Less Sycophantic validation accuracy up to 16% respectively compared to reward models trained on our main datasets. Interestingly, the validation accuracy of Concise and Less Sycophantic increases beyond the validation accuracy of reward models trained on clean datasets when oversampling by $\alpha=150$%. We hypothesize that flipping negatively-contributing samples not only corrects manually flipped preference labels but also helps remove pre-existing biased samples in the Anthropic-HH dataset. Meanwhile, the validation accuracy of Verbose and More Sycophanctic slightly decreases compared to the reward model trained using our main dataset, indicating a tradeoff between reducing bias and maintaining performance on these sets. However, the magnitude of this decrease is relatively small compared to the substantial improvements observed for the Concise and Less Sycophantic sets.
Length Bias
Dataset Concise Val Acc. (%) Verbose Val Acc. (%)
Length (main) 49.93 44.93
+ Flipped ($3.28$%) 54.54 (+4.61) 42.84 (-2.09)
+ Flipped ($6.56$%) 59.37 (+9.44) 40.78 (-4.15)
+ Flipped ($9.84$%) 64.12 (+14.20) 38.07 (-6.86)
Clean Dataset 58.53 (+8.61) 41.82 (-3.11)
Sycophancy Bias
Train Dataset Less Syco. Val Acc. (%) More Syco. Val Acc. (%)
Sycophancy (main) 55.56% 74.00
+ Flipped ($2.09$%) 61.99% (+6.43) 73.33 (-0.67)
+ Flipped ($4.17$%) 61.99% (+6.43) 69.33 (-4.67)
+ Flipped ($6.26$%) 71.93% (+16.37) 68.00 (-6.00)
Clean Dataset 62.57% (+7.02) 74.00 (+0.00)
Table 10: Comparison of reward model validation accuracy after curating the training dataset. "Flipped ($r$%)" indicates that $r$of the dataset were flipped. Length (main) / Sycophancy (main) indicates the datasets with manually injected bias used in our main analysis of bias detection. Clean dataset indicates a dataset without manually injected bias. The symbol ($\pm$%) indicates the difference to Length Original and Sycophancy Original.
Appendix H Qualitative Analysis
@@
To investigate length bias, we used the Concise dataset to calculate influence values, focusing on cases where response length may affect outcomes. As shown in Figure 9, the most positively-contributing samples, which negatively impact bias, show no flipped instances, and the selected responses demonstrate no preference for length. This indicates that these responses did not favor verbose answers.
@@
For sycophancy bias, we examined the Less Sycophantic dataset to calculate influence values. As illustrated in Figure 10, most positively-contributing samples do not contain any flipped instances, and none of the responses show sycophancy bias. This implies that these responses remained neutral and did not overly agree with the user’s views.
@@
Figure 9: The most positively-contributing samples of the length bias experiment do not contain any flipped instances, and the selected responses show no evidence of length bias. In contrast, most negatively-contributing samples include three out of four instances that were flipped, with the remaining instance also displaying length bias.
Figure 10: The most positively-contributing samples of the sycophancy bias experiment do not include any flipped instances, and the selected responses show no signs of sycophancy bias. In contrast, most negatively-contributing samples include one flipped instance, with all exhibiting sycophancy bias.
Appendix I Ablation Experiments on Bias Detection
In this section we study the effects of validation set composition in Section I.1 and validation set size in Section I.2, on the performance of bias detection using influence functions. We also study the effect of increasing the number of few-shot examples for LLM baselines in Section I.3. Additionally, we investigate how performance is affected when validation and training sets differ in distribution in Section I.4.
I.1 Validation Set Ablation
I.2 Validation Set Size Ablation for Influence Functions
@@
The ablation results of the validation set size are given in Figure 12. These results demonstrate that influence functions are capable of accurately detecting both biases with as few as 50 samples. Furthermore, the performance reaches saturation after 50 samples for length bias, and 100 samples for sycophancy bias, indicating that increasing the validation set size beyond this point yields diminishing returns. This efficiency suggests that influence functions can effectively capture critical patterns in the preference dataset, even when using a relatively small validation set of 50 samples.
I.3 Few-Shot Example Ablation for LLM Baselines
@@
I.4 Validation Distribution Ablation using HelpSteer2 Validation Set
Training Set Source Validation Set Source AUC
Anthropic-HH Anthropic-HH 0.800
Anthropic-HH HelpSteer2 0.620
Table 11: Impact of validation set distribution on AUC performance of influence functions. We observe a drop from 0.800 to 0.620.
To investigate how distributional differences impact performance, we retain the Anthropic-HH Bai et al. (2022) training set but substitute the validation set with samples from the HelpSteer2 Wang et al. (2024) dataset. Specifically, we conduct an additional experiment using our length bias detection setup. For this new validation set, we select responses from the existing HelpSteer2 validation split by defining responses with higher helpfulness scores as ’winning’ and those with lower scores as ’losing,’ discarding any ties. We then specifically focus on cases where the winning response is shorter, resulting in a final validation set of 145 samples.
We emphasize the significant distributional differences between these two datasets. Prompts in HelpSteer2, sourced from the crowd-sourced ShareGPT RyokoAI (2023) dataset, encompass diverse real-world interactions between real users and ChatGPT, ranging from brief keyword-based inputs (e.g., ‘c#’) to complex, multi-step instructions. In contrast, Anthropic-HH prompts typically feature straightforward, explicitly structured Q&A interactions, such as ‘What’s the best way to travel from NYC to Boston?’ Additionally, responses in HelpSteer2 were generated using the more recent Nemotron 4 model Adler et al. (2024), whereas Anthropic-HH utilized an earlier-generation model.
Despite significant distributional differences between the training set and validation set, the influence function surpassed baselines like Confidence, Mahalanobis. And Gemini-1.5-Pro. Specifically, using the HelpSteer2 validation set resulted in an AUC of 0.620, lower than the original 0.800 obtained with the Anthropic-HH Concise set. This indicates that aligning the validation set distribution closely with the training set is important to achieve optimal performance. Nevertheless, results indicate that influence functions remain effective for detecting biased samples even when applied across differing distributions.
Appendix J Detecting Biased Samples in the Original Human Datatset
We conduct a human survey to evaluate whether influence function values, estimated using the Concise and Less Sycophantic sets, effectively detect length-biased and sycophancy-biased preference labels in the original Anthropic-HH dataset. Our experiments use the same training dataset as our main study but retain the original preference labels. All other training details remain consistent with our bias detection experiments. We sample the top 100 negatively-contributing samples (Top-100) using influence functions and a randomly selected set of 100 samples (Random-100) for comparison. Note that the Random-100 subsets for the length bias and sycophancy bias experiments are drawn from their corresponding training datasets of length and sycophancy. Human annotators then evaluate the responses in each set, selecting more helpful responses. We then compare these annotations with the original Anthropic-HH labels to determine whether mislabeled preference samples occur more frequently in the Top-100 set.
Annotation Procedure
We recruit four human annotators (two authors and two non-authors) to evaluate response helpfulness based on the conversation. Annotators were carefully instructed to fully understand the conversations before proceeding with annotation. While the authors participated as annotators, all annotators remained unaware of the original preference labels, as the response pairs were anonymized and shuffled. Following the dataset collection process of Anthropic-HH (Bai et al., 2022), annotators are instructed to ‘Choose the most helpful and honest response’. Since many samples in Anthropic-HH contain responses of similar quality, we also allow for a tie when no clear preference is available. The response pairs of the samples are anonymized and shuffled to ensure that the annotators are not aware of the original preference label of the Anthropic-HH dataset.
Results
Table 12 shows that in the length bias experiment, 47 samples were mislabeled in Top-100 compared to only 13 samples in Random-100. In the sycophancy bias experiment, 55 samples were mislabeled in Top-100 compared to 33 samples in Random-100. These results demonstrate the effectiveness of our method in detecting labeler bias in real datasets. Furthermore, they suggest that such biases are not uncommon in real datasets like Anthropic-HH, further highlighting the need for accurate bias detection methods like ours.
Length Bias
Subset Mislabeled Correct Tie
Top-100 47 38 15
Random-100 13 69 18
Sycophancy Bias
Subset Mislabeled Correct Tie
Top-100 55 36 9
Random-100 33 59 8
Table 12: Human survey results comparing Top-100 subsets selected using our influence function approach with Random-100 subsets for length and sycophancy experiments. Significant portions of Top-100 samples are mislabeled compared to Random-100 for both biases.
Examples of mislabeled samples
We provide examples of mislabeled samples in Figure 15 and Figure 16, which demonstrates two key patterns: (1) factually accurate and appropriate responses being rejected in favor of verbose or misleading alternatives, and (2) overly sycophantic or subjective responses being preferred over more objective ones. Preference between responses may vary depending on the human annotator and their interpretation of what is ‘helpful’. However, we observe a significant quality difference between response pairs in the negatively-contributing samples selected using influence functions. In the length bias experiments, the chosen responses in the Top-100 set often contained factually inaccurate information, whereas the rejected responses were relatively more accurate. In the sycophancy bias experiment, chosen responses of the Top-100 set included excessive apologies, personal tone, or subjective statements compared to fact-based and more neutral rejected responses. While a degree of agreement and politeness can enhance helpfulness, such responses should be grounded in factual accuracy and objectivity. These findings suggest labeler biases in human annotations are not uncommon in real-world data, emphasizing the need for accurate labeler bias detection methods like ours to improve the reliability of preference datasets.
Appendix K Ablation Experiments on Labeling Strategy Oversight
We measure the Label Accuracy and Cosine Similarity metrics when estimating influence functions on different validation set sizes for our labeler strategy oversight experiment. Specifically, we random sample 20 subsets from the full validation set, and average the metrics. Figure 14 shows that our approach performs reasonably well even with 50 validation samples. This demonstrates the efficiency of our method, even in scenarios where obtaining expert-labeled data is challenging.
Figure 15: Examples of the Top 100 negatively-contributing samples measured by the Concise set. In all three examples, the originally rejected response is more factually accurate or appropriate. To enhance clarity, we use color coding: green for factually correct and appropriate statements, orange for mildly misleading or inaccurate information, and red for severely incorrect or misleading content.
Figure 16: Examples of the Top 100 negatively-contributing samples measured by the Less Sycophantic set. In all examples, the originally chosen response exhibits excessive subjectivity or flattery. To improve clarity, we use color coding: orange for sycophantic content and red for misleading information.